y = x + x-2 + ... + 1, where n + 1. where n = odd and f(5) = 137. Find the value of f(-5). Done
Answer #1

+ x ny + y = x²+x-2 +1 = f(x) (x²+x2-2 + ) = f(a)-1 for x=5 (5^+50-2 ...) = f(5)-1 137-1 = 136 for x=-5 ((-5) + (-5.jon 4.-)